OpenStudy (anonymous):

one more question thanks! solve formula for n: s=(n-2)180

OpenStudy (saifoo.khan):

Divide 180 from both sides,\[\frac{s}{180} = n - 2\]Add 2 on both sides, \[n = 2 + \frac{s}{180}\]

OpenStudy (anonymous):

n=2+s180

jhonyy9 (jhonyy9):

s=180n-360 s+360=180n (s+360)/180 =n